# Cosmic Web Classification through Stochastic Topological RankingCosmic Web Classification through Stochastic Topological Ranking

### v2

@@ 2 Methodology
Figure 1: Illustration of the ASTRA method applied to a 2-dimensional dataset of 50 points. The left panel displays the input object points (black points) and the randomly generated points uniformly distributed in the Monte Carlo iteration (blue points, smaller than the object points). The right panel depicts the Delaunay graph, with solid lines connecting two object points and dashed lines connecting at least one random point. In addition, this graph has 2 zoom-ins (of areas with high and low density from left to right), which qualitatively show the difference between the proportion of the number of connections of the object points between them over the connections with random points. In the HD region solid lines (connections to object points) dominate, while in the LD region dashed lines (connections to random points) dominate.
@@ 2.1 Algorithm Description
    =$\mathcal{R}=\{r_{1},r_{2},...,r_{N_{R}}\}$: The set of randomly distributed points that follow the same selection function and geometrical constraints as $\mathcal{O}$. We refer to this as the "random catalog".
In our implementation, we generate an equal number of random points as objects, i.e., $N_{R}=N_{O}$, to ensure consistent local mean point density in both sets. This also necessitates that the volume spanned by $\mathcal{O}$ and $\mathcal{R}$ remains the same.
The merged catalog $\mathcal{M}=\mathcal{O}\cup\mathcal{R}$, comprising both the object and random catalogs, undergoes Delaunay triangulation. For each point $p_{i}\in\mathcal{M}$, we compute:
@@ 2.3 The importance of $N_{R}=N_{O}$
The choice of $\rho=1$ is particularly effective for identifying random points in underdense regions (voids) that are typically challenging to characterize due to the lack of galaxies in these areas. The equal number of random points ensures these regions are well-sampled. This aspect of ASTRA is especially valuable, as it addresses a common limitation in cosmic web classification methods. Furthermore, as showed in Gómez Cortés (2019), using values of $\rho>1.6$ results in void percolation across the volume, i.e., identified voids start to increase in size due to a percolation process that links up voids across the volume. For large values of $\rho\approx 2$ the percolation tipically reaches a state where a single void has merged into itself all of the random points classified as voids.
@@ 2.4 Threshold Determination
The classification thresholds in Table 1 are derived through a statistical procedure designed to ensure the method’s efficacy across varying density distributions. The process for determining these thresholds is as follows.
We generate two sets of points with identical spatial extents and number densities. The first set represents a homogeneous distribution mimicking the observed points, while the second set replicates the characteristics of the random points. This construction maintains the $\rho=1$ ratio between random and data points, as previously discussed.
Using these generated point sets, we apply the ASTRA methodology to compute the $r$ values for each point. The resulting distribution of $r$ values is then analyzed to determine the upper and lower thresholds that encompass 99% of the points around the mean of $r=0$. This procedure yields the threshold values of $-0.9$ and $0.9$ for the lower and upper bounds, respectively.
The choice of these thresholds has the following interpretation: in a statistically homogeneous number density distribution, we expect no detection of extreme density regions (voids or knots). By setting the thresholds at the 99th percentile, we ensure that ASTRA classifies only the most statistically significant deviations from homogeneity as voids or knots. This approach strikes a balance between sensitivity to genuine cosmic web structures and robustness against statistical fluctuations. Finally, the intermediate threshold between sheets and filaments is simply determined by the underdense/overdense criterion around $r=0$.
It is worth noting that while these thresholds are derived from idealized homogeneous distributions, they prove effective in classifying real cosmic web structures due to the algorithm’s local nature and the stochastic properties of the random point distribution. The inclusion of random points in the classification process, the key innovation of ASTRA, allows for accurate identification of underdense regions (voids) that are typically challenging to characterize due to the lack of galaxies in these areas.
@@ 3 Data
We use three different datasets as input to ASTRA. The first simulated dataset used in this study is from the Tracing the Cosmic Web (TCW) comparison project (Libeskind et al., 2017). The simulation is a dark matter-only N-body simulation, with a box size of 200h-1Mpc and 5123 particles, using the Gadget-2 code (Springel, 2005) and cosmological parameters of $h=0.68$, $\Omega_{M}=0.31$, $\Omega_{\Lambda}=0.69$, $n_{s}=0.96$, and $\sigma_{8}=0.82$. A dataset of 281,465 dark matter haloes was obtained using a Friends-of-Friends (FOF) algorithm (Davis et al., 1985) with a linking length of $b=0.2$ and a minimum of 20 particles for this analysis. This simulation is advantageous as it has already been analyzed by 11 different methods, which have also classified each point into one of the four cosmic structures.
@@
The third catalog used in our study is an observational data catalog from the Sloan Digital Sky Survey (SDSS) Data Release 7 (Abazajian et al., 2009). Specifically, we use data from the NYU Value-Added Galaxy Catalog (Blanton et al., 2005), which includes large-scale structure samples constructed from SDSS data. Initially, we had 559,028 galaxies. To create a volume-limited sample, we applied cuts to the r-band magnitude and redshift. We selected all galaxies with an r-band magnitude of $M_{r}\leq-20$ and a redshift $z\leq 0.114$. We chose this magnitude limit to preserve all galaxies brighter than $L_{*}$ (Pan et al., 2012), as they reveal the structure of the cosmic web. Additionally, we selected galaxies in the declination range between 0° and 50°, and right ascension range between 120° and 230°, resulting in a final sample of 90,655 galaxies. This specific angular range was chosen to simplify the process of generating random points. We compute the Cartesian coordinates through a two-step process. First, we convert the redshift into a comoving radial distance using the cosmological parameters from the TNG simulation. Then, we utilize the angular coordinates (right ascension and declination) along with the radial coordinate to compute the final $x$, $y$, $z$ Cartesian coordinates. It is important to note that we do not apply corrections for redshift space distortions. This limitation may lead to an overclassification of galaxies as filaments, as high-density peaks are elongated along the line of sight in redshift space (Fang et al., 2019), potentially blending into neighboring filamentary regions. It is important to note that neither of the simulation-based catalogs (TCW and TNG) include modeling of observational effects such as survey incompleteness or redshift space distortions.
@@ 4 Results
Figure 3: Plot of the classification given by ASTRA for the uniformly generated random points in each of the three data catalogs. following the same pattern as in Figure 2. The structure of each of the painted points was chosen with the results of the iteration classification. The most relevant characteristing is that now voids are clearly visible by multiple random tracers.
@@ 4.1 Visual inspection
Since cosmic web algorithms are designed and tuned to various degrees to reproduce visual impressions (Libeskind et al., 2017), we begin with a qualitative exploration of these visual patterns before proceeding to quantitative analysis.
@@ 4.3 Volume Filling Fractions and Mass Segregation
Comparing these results with other methods from the TCW paper (Libeskind et al., 2017), we find notable differences in VFF values. As shown in Figure 5 and Table 2 of Libeskind et al. (2017), most cosmic web classification methods report void VFF values ranging from 40% to 70%. For instance, NEXUS identifies void volumes of 65% and T-web reports 43%. DisPerSE, the only other method in TCW that classifies cosmic web types from sparse data, reports VFF values of 23.9%, 37.3%, and 38.8% for filaments, sheets, and voids, respectively. ASTRA finds substantially smaller void volumes (7% versus 38.8%) and larger sheet volumes (70% versus 37.3%).
@@ 4.4 Void Catalogs
The viability of ASTRA’s voids for statistical applications is supported by several key results: (1) the void size function follows theoretical expectations (Figure 9), (2) strong correlation signals in void-galaxy cross-correlations (Figure 10), (3) mass segregation patterns showing the expected lowest-mass galaxies in voids (Figure 7), and (4) visual inspection confirming these regions correspond to genuine underdense areas (Figures 2 and 3). These results demonstrate that while ASTRA identifies smaller void volumes, these regions represent physically meaningful void cores suitable for statistical analysis.
@@
In our approach, we utilize random points that sample voids, not just their centers. With the number of random points matching the number of galaxies, and approximately 10% of random points tracing voids, the galaxy sample of $10^{6}$ can be cross-correlated with $10^{5}$ positions representing voids. This increase of two orders of magnitude in the number of galaxy-void pairs used for the 2D cross-correlation function enhances precision. While a thorough examination of this method’s potential for constraining cosmological parameters is reserved for future work, we explore its feasibility by analyzing the results of measuring the auto and cross-correlation functions for different web types in both data and random samples, as measured by ASTRA.
@@
The main results for these correlations are shown in Figure 10. The monopole exhibits its largest amplitudes for the auto-correlation of data-filaments, random-voids, and their cross-correlation. Furthermore, these correlations show a distinctive transitional scale around $20$ Mpc. The next high-amplitude cross-correlation is found for data-sheets and random-voids, also with a transitional scale around $15$-$20$ Mpc. Some of the quadrupoles show large anisotropies, the largest being found for the auto-correlation of data-filaments and their cross-correlation with random-voids, especially for scales larger than $40$ Mpc. This might be due to the fact that filaments and voids are expected to show large redshift space distortions which influence their selection effect when identified in redshift space.
@@
Table 3: Overview of the methods used in the Tracing the Cosmic Web (Libeskind et al., 2017). The Input distinguishes between dense (tipycally DM computational particles from an N-body simulation) and sparse (typically DM haloes or galaxies).
@@ 4.6 Comparison Against Other Cosmic Web Finding Methods
Table 3 provides an overview of the methods employed in the TCW project, for which public data exists on the classification of FOF DM haloes. The table outlines the web types each method can categorize, the types of input data it can handle (dense, from simulations, or sparse, from observations), and whether it operates on a grid-based system. To assess the similarity between the results obtained using different methods and those obtained using ASTRA, we computed confusion matrices and calculated the weighted average F1 score across different structures.
Figure 11 presents the confusion matrices for the 11 methods compared to the results of all methods in the TCW paper (Libeskind et al., 2017). Two key observations emerge from this figure. Firstly, the most significant discrepancies in classification occur for voids, where ASTRA often classifies DM halos as sheets, contrary to other methods. Secondly, the highest level of agreement in classification generally occurs for filaments. The classification accuracy across the six methods that categorize into four structures typically follows a decreasing order: filaments, sheets, knots, and voids.
To provide a quantitative evaluation of these observations, we employed the F1 score, a commonly used metric in statistical analysis to assess classification accuracy. The F1 score is calculated as the harmonic mean of precision and recall, where precision denotes the ratio of true positive results to the total number of positive results found, and recall represents the ratio of true positive results to the total number of results that should have been classified as positive. For methods classifying three or more classes, we utilized the F1 score weighted by the number of instances in each class.
@@ 5 Conclusions
Our application of ASTRA to three distinct datasets—a dark matter-only simulation, a hydrodynamical simulation, and SDSS observational data—yields several insights:
@@
While ASTRA shows promise, we acknowledge several limitations that need attention in future work. The method currently uses conservative threshold values ($r\leq-0.9$ for voids) that identify only the most extreme void cores, differing from broader void definitions in some existing methods. However, this represents a design choice rather than a limitation—the $r$ parameter framework allows users to adjust thresholds for their specific research questions.
For instance, relaxing the void threshold would increase void volume filling fractions to match some literature values, while different filament/sheet boundaries could highlight various structural features.
@@
Despite these limitations, our comparative analysis with other cosmic web finding techniques shows that ASTRA provides classifications that align with established methods, particularly for filamentary structures. The method’s main strengths—working with sparse data, handling both overdense and underdense regions, and avoiding smoothing or interpolation—make it a useful addition to existing cosmic web analysis tools.
Going forward, we expect that ASTRA’s approach to cosmic web classification, particularly its identification of various environment types and use of random points for underdense region mapping, will be useful for current and upcoming large spectroscopic surveys.
The method’s ability to create detailed void catalogs and measure cross-correlations between different cosmic web elements may improve cosmological parameter estimation and expand our understanding of structure formation in the Universe.

### v3

@@ 2 Methodology
Figure 1: Illustration of the ASTRA method applied to a 2-dimensional dataset of 50 points. The left panel displays the input object points (larger black points) and the randomly generated points uniformly distributed in the Monte Carlo iteration (blue points, smaller than the object points). The right panel depicts the Delaunay graph, with solid lines connecting two object points and dashed lines connecting at least one random point. In addition, this graph has 2 zoom-ins (of areas with high and low density from left to right), which qualitatively show the difference between the proportion of the number of connections of the object points between them over the connections with random points. In the HD region solid lines (connections to object points) dominate, while in the LD region dashed lines (connections to random points) dominate.
@@ 2.1 Algorithm Description
    =$\mathcal{R}=\{r_{1},r_{2},...,r_{N_{R}}\}$: The set of randomly distributed points that follow the same selection function and geometric constraints as $\mathcal{O}$. We refer to this as the "random catalog". It represents an unclustered Poisson sampling of a constant background density.
The random catalog follows the same selection function and geometrical constraints as the object catalog, meaning it mimics the survey’s observational limitations and data processing cuts. For observational data, this includes reproducing the angular selection boundaries (such as the declination and right ascension ranges), the radial distance distribution derived from the redshift selection cuts, and any magnitude or completeness limits applied to the galaxy sample. For simulation data, the random points are uniformly distributed within the same volume boundaries as the simulated objects. This approach follows standard practices in large-scale structure analysis, where random catalogs are essential for accurate correlation function measurements and statistical analyzes (Landy & Szalay, 1993; Hamilton, 1993).
In our implementation, we generate an equal number of random points as objects, i.e. $N_{R}=N_{O}$, to ensure a consistent local mean point density in both sets. This also requires that the volume spanned by $\mathcal{O}$ and $\mathcal{R}$ remains the same.
The merged catalog $\mathcal{M}=\mathcal{O}\cup\mathcal{R}$, comprising both the object and the random catalogs, undergoes Delaunay triangulation. For each point $p_{i}\in\mathcal{M}$, we compute the following:
@@
Figure 1 illustrates the ASTRA methodology applied to a two-dimensional example with 50 points. The left panel shows the input object points (black) and randomly generated points (blue, smaller), while the right panel displays the resulting Delaunay graph with connections between points. The zoomed insets demonstrate how the proportion of connections varies between high-density (HD) regions (where solid lines connecting object points dominate) and low-density (LD) regions (where dashed lines connecting to random points are more prevalent). This connectivity pattern forms the basis for our cosmic web classification scheme.
@@ 2.2 Choice of Graph Method
Furthermore, the use of Delaunay tessellations for cosmic web analysis has a rich history in the field. The Delaunay Tessellation Field Estimator (DTFE) developed by Schaap & van de Weygaert (2000); van de Weygaert & Schaap (2009) established the foundation for exploiting the geometrically adaptive properties of Delaunay tessellations in cosmological applications. DTFE has demonstrated exceptional performance in reconstructing density and velocity fields from discrete point distributions while preserving both the multiscale character and local geometry of cosmic structures (Cautun & van de Weygaert, 2011).
The DTFE methodology has subsequently been incorporated into several cosmic web classification algorithms. The Multiscale Morphology Filter (MMF) and its successor NEXUS (Aragón-Calvo et al., 2007; Cautun et al., 2012) utilize DTFE-reconstructed density fields as the basis for their scale-space analysis of cosmic web morphology. Similarly, the Watershed Void Finder (WVF) (Platen et al., 2007) employs DTFE density reconstructions in conjunction with watershed techniques for void identification. The topological analysis framework DisPerSE also builds upon DTFE density field reconstructions (Sousbie, 2011).
While ASTRA shares with these methods the fundamental insight that Delaunay tessellations provide an optimal framework for cosmic web analysis, our approach differs in its direct use of the tessellation connectivity without requiring density field reconstruction or interpolation. Rather than first estimating a continuous density field and then applying morphological filters, ASTRA operates directly on the discrete connectivity patterns of the Delaunay graph, combined with the novel inclusion of random points to characterize both overdense and underdense regions.
@@ 2.3 The importance of $N_{R}=N_{O}$
The choice of $\rho=1$ is particularly effective in identifying random points in underdense regions (voids) that are typically challenging to characterize due to the lack of galaxies in these areas. The equal number of random points ensures that these regions are well-sampled. This aspect of ASTRA is especially valuable, as it addresses a common limitation in cosmic web classification methods. Furthermore, as shown in Gómez Cortés (2019), using values of $\rho>1.6$ results in percolation of voids throughout the volume, that is, identified voids start to increase in size due to a percolation process that links the voids throughout the volume. For large values of $\rho\approx 2$ the percolation typically reaches a state where a single void has merged into itself all the random points classified as voids.
@@ 2.4 Threshold Determination
The classification thresholds listed in Table 1 are established through a statistical procedure designed to maintain robustness across varying density environments. The process for determining these thresholds is as follows.
Two synthetic point sets are generated within identical spatial boundaries and with equal number densities. One set emulates a homogeneous distribution, analogous to the observed data, while the other replicates the properties of the accompanying random catalog. This setup preserves the $\rho=1$ ratio between data and random points, as previously discussed.
The ASTRA algorithm is then applied to both sets, computing $r$ values for each point. The distribution of these $r$ values is analyzed to identify symmetric thresholds that encompass 99% of the data centered on $r=0$. This yields cutoff values of 0.9-0.9- 0.9 and 0.90.90.9, corresponding to the lower and upper bounds, respectively.
[…]
In the context of a statistically uniform distribution, these thresholds imply that voids and knots should not be detected, ensuring that ASTRA highlights only the most statistically significant fluctuations. This criterion effectively balances sensitivity to meaningful cosmic web structures with resistance to noise. The intermediate transition between sheets and filaments naturally arises from the sign of $r$, distinguishing underdense from overdense environments around $r=0$.
Empirical tests confirm that the 99th percentile threshold yields classifications with appropriate mass segregation and volume-filling fractions. While somewhat subjective, like threshold choices in T-web, V-web, DisPerSE, or watershed-based methods, this selection performs well across a range of applications. Users may tailor the threshold to suit their scientific goals.
Although derived from idealized, homogeneous configurations, these thresholds remain effective for real data due to ASTRA’s local approach and the inherent randomness of the point distribution. The inclusion of random points, the key innovation of ASTRA, enables the detection of voids, which are otherwise difficult to characterize due to their sparsity in galaxy surveys.
@@ 3 Data
We use three different datasets as input to test ASTRA. The first simulated dataset used in this study is from the Tracing the Cosmic Web (TCW) comparison project (Libeskind et al., 2018). The simulation is a dark matter-only N-body simulation, with a box size of 200 Mpc $h^{-1}$ and 5123 particles, using the Gadget-2 code (Springel, 2005) and cosmological parameters of $h=0.68$, $\Omega_{M}=0.31$, $\Omega_{\Lambda}=0.69$, $n_{s}=0.96$, and $\sigma_{8}=0.82$. A dataset of 281,465 dark matter haloes was obtained using a Friends-of-Friends (FOF) algorithm (Davis et al., 1985) with a linking length of $b=0.2$ and a minimum of 20 particles for this analysis, resulting in a number density of $11.0\times 10^{-3}$ Mpc-3. This simulation is advantageous as it has already been analyzed by 11 different methods, which have also classified each point into one of the four cosmic structures.
@@
The third catalog used in our study is an observational data catalog from the Sloan Digital Sky Survey (SDSS) Data Release 7 (Abazajian et al., 2009). Specifically, we use data from the NYU Value-Added Galaxy Catalog (Blanton et al., 2005), which includes large-scale structure samples constructed from SDSS data. Initially, we had 559,028 galaxies. To create a volume-limited sample, we applied cuts to the r-band magnitude and redshift. We selected all galaxies with an r-band magnitude of $M_{r}\leq-20$ and a redshift $z\leq 0.114$. We chose this magnitude limit to preserve all galaxies brighter than $L_{*}$ (Pan et al., 2012), as they reveal the structure of the cosmic web. Additionally, we selected galaxies in the declination range between 0° and 50°, and right ascension range between 120° and 230°, resulting in a final sample of 90,655 galaxies. This corresponds to a total comoving volume of $5.2\times 10^{7}$ Mpc3 and a number density of $1.7\times 10^{-3}$ Mpc-3 This specific angular range was chosen to simplify the process of generating random points. Then, we compute the Cartesian coordinates through a two-step process. First, we convert the redshift into a comoving radial distance using the cosmological parameters from the TNG simulation. Then, we utilize the angular coordinates (right ascension and declination) along with the radial coordinate to compute the final $x$, $y$, $z$ Cartesian coordinates. It is important to note that we do not apply corrections for redshift space distortions. This limitation may lead to an overclassification of galaxies as filaments, as high-density peaks are elongated along the line of sight in redshift space (Fang et al., 2019), potentially blending into neighboring filamentary regions. It is important to note that neither of the simulation-based catalogs (TCW and TNG) include modeling of observational effects such as survey incompleteness or redshift space distortions.
It is important to acknowledge that our three datasets represent a heterogeneous collection with fundamentally different selection criteria and object types. The TCW catalog contains dark matter halos with a minimum mass threshold of approximately $2.0\times 10^{12}M_{\odot}$ $h^{-1}$, the TNG catalog includes galaxies selected by stellar mass above $10^{9}M_{\odot}$ $h^{-1}$, and the SDSS sample comprises galaxies selected by r-band luminosity ($M_{r}\leq-20$). These different selection functions result in varying number densities: TCW ($11.0\times 10^{-3}$ Mpc-3), TNG300 ($8.2\times 10^{-3}$ Mpc-3), and SDSS ($1.7\times 10^{-3}$ Mpc-3). While this heterogeneity demonstrates ASTRA’s versatility in handling diverse data types, from dark matter halos to observed galaxies, it does introduce limitations in making direct quantitative comparisons between datasets. The different mass/luminosity thresholds and sampling strategies mean that each catalog traces different populations of objects and potentially different aspects of the underlying cosmic web structure. Nevertheless, the broad similarity in number densities between TNG and TCW, combined with the ability to analyze sparse observational data like SDSS, showcases the method’s applicability across the range of datasets commonly used in large-scale structure studies.
@@ 4 Results
Figure 3: Plot of the classification given by ASTRA for the uniformly generated random points in each of the three data catalogs. following the same pattern as in Figure 2. The structure of each of the painted points was chosen with the results of the iteration classification. The most relevant characteristic is that now voids are clearly visible by multiple random tracers.
@@ 4.1 Visual inspection
Since cosmic web algorithms are designed and tuned to various degrees to reproduce visual impressions (Libeskind et al., 2018), we begin with a qualitative exploration of these visual patterns before proceeding to quantitative analysis.
@@ 4.3 Volume Filling Fractions and Mass Segregation
Comparing these results with other methods from the TCW paper (Libeskind et al., 2018), we find notable differences in VFF values. As shown in Figure 5 and Table 2 of Libeskind et al. (2018), most cosmic web classification methods report void VFF values ranging from 40% to 70%. For instance, NEXUS identifies void volumes of 65% and T-web reports 43%. DisPerSE, the only other method in TCW that classifies cosmic web types from sparse data, reports VFF values of 23.9%, 37.3%, and 38.8% for filaments, sheets, and voids, respectively. ASTRA finds substantially smaller void volumes (7% versus 38.8%) and larger sheet volumes (70% versus 37.3%).
@@ 4.4 Void Catalogs
The viability of ASTRA’s voids for statistical applications is supported by several key results: (1) visual inspection confirming these regions correspond to genuine underdense areas (Figures 2 and 3), (2) mass segregation patterns showing the expected lowest-mass galaxies in voids (Figure 7), and (3) the void size function follows theoretical expectations (Figure 9). These results demonstrate that while ASTRA identifies smaller void volumes, these regions represent physically meaningful void cores suitable for statistical analysis.
@@
In our approach, we utilize random points that sample voids, not just their centers. With the number of random points matching the number of galaxies, and approximately 10% of random points tracing voids, the galaxy sample of $10^{6}$ can be cross-correlated with $10^{5}$ positions representing voids. This increase of two orders of magnitude in the number of galaxy-void pairs used for the 2D cross-correlation function enhances precision. While a thorough examination of this method’s potential for constraining cosmological parameters is reserved for future work, we explore in the next section its feasibility by analyzing the results of measuring the auto and cross-correlation functions for different web types in both data and random samples, as measured by ASTRA.
@@
The main results for these correlations are shown in Figure 10. The monopole exhibits its largest amplitudes for the auto-correlation of data-filaments, random-voids, and their cross-correlation. Furthermore, these correlations show a distinctive transitional scale around 202020 Mpc. The next high-amplitude cross-correlation is found for data-sheets and random-voids, also with a transitional scale around 151515-202020 Mpc. Some of the quadrupoles show large anisotropies, the largest being found for the auto-correlation of data-filaments and their cross-correlation with random-voids, especially for scales larger than 404040 Mpc. This might be due to the fact that filaments and voids are expected to show large redshift space distortions which influence their selection effect when identified in redshift space.
@@
Table 3: Overview of the methods used in the Tracing the Cosmic Web (Libeskind et al., 2018). The Input distinguishes between dense (tipycally DM computational particles from an N-body simulation) and sparse (typically DM haloes or galaxies).
@@ 4.6 Comparison Against Other Cosmic Web Finding Methods
Table 3 provides an overview of the methods employed in the TCW project, for which public data exists on the classification of FOF DM haloes. The table outlines the web types each method can categorize, the types of input data it can handle (dense, from simulations, or sparse, from observations), and whether it operates on a grid-based system. To assess the similarity between the results obtained using different methods and those obtained using ASTRA, we computed confusion matrices and calculated the weighted average F1 score across different structures. The F1 score is calculated as the harmonic mean of precision and recall, where precision denotes the ratio of true positive results to the total number of positive results found, and recall represents the ratio of true positive results to the total number of results that should have been classified as positive.
Figure 11 presents the confusion matrices for the 11 methods compared to the results of all methods in the TCW paper (Libeskind et al., 2018). Two key observations emerge from this figure. Firstly, the most significant discrepancies in classification occur for voids, where ASTRA often classifies DM halos as sheets, contrary to other methods. Secondly, the highest level of agreement in classification generally occurs for filaments. The classification accuracy across the six methods that categorize into four structures typically follows a decreasing order: filaments, sheets, knots, and voids.
@@ 5 Conclusions
Our application of ASTRA to three distinct datasets (a dark matter-only simulation, a hydrodynamical simulation, and SDSS observational data) yields several insights:
@@
While ASTRA shows promise, we acknowledge several limitations that need attention in future work. The method currently uses conservative threshold values ($r\leq-0.9$ for voids) that identify only the most extreme void cores, differing from broader void definitions in some existing methods. However, this represents a design choice rather than a limitation: the $r$ parameter framework allows users to adjust thresholds for their specific research questions. For instance, relaxing the void threshold would increase void volume filling fractions to match some literature values, while different filament/sheet boundaries could highlight various structural features.
@@
Despite these limitations, our comparative analysis with other cosmic web finding techniques shows that ASTRA provides classifications that align with established methods, particularly for filamentary structures. The method’s main strengths (working with sparse data, handling both overdense and underdense regions, and avoiding smoothing or interpolation) make it a useful addition to existing cosmic web analysis tools. The method’s ability to create detailed void catalogs and measure cross-correlations between different cosmic web elements may improve cosmological parameter estimation and expand our understanding of structure formation in the Universe.
Going forward, we expect that ASTRA’s approach to cosmic web classification, particularly its identification of various environment types and use of random points for underdense region mapping, will be useful for current and upcoming large spectroscopic surveys. In particular, we plan to apply our algorithm to the data from the Dark Energy Spectroscopic Instrument (DESI), taking particular interest into the Large Scale Structure catalogs provided by the collaboration (Ross et al., 2025), which already include carefully constructed randoms as required by ASTRA.
ACKNOWLEDGEMENTS
We thank our reviewers, including Miguel Aragon-Calvo, for their valuable insights and specific advice that greatly improved this manuscript.
JEFR acknowledges funding from Facultad de Ciencias at Universidad de los Andes under project INV-2023-162-2846.
CONFLICT OF INTEREST
Authors declare no conflict of interest.